OpenStudy (anonymous):

Given the following system of equations, identify the type of system. y = -x + 4 2x + 2y = 8 A. inconsistent b. equivalent

jimthompson5910 (jim_thompson5910):

are you able to solve 2x + 2y = 8 for y?

OpenStudy (anonymous):

i dont know how. van you explain?

jimthompson5910 (jim_thompson5910):

think of 2x+2y as 2y+2x what operation will undo that +2x ?

OpenStudy (anonymous):

subtract 2 from both sides?

jimthompson5910 (jim_thompson5910):

2x actually

jimthompson5910 (jim_thompson5910):

so we will have 2y = 8 - 2x which is the same as 2y = -2x + 8

jimthompson5910 (jim_thompson5910):

we want to isolate y fully so we have to move that 2 (from 2y) over to the other side. How do we do this?

jimthompson5910 (jim_thompson5910):

hint: 2y means "2 times y"

OpenStudy (anonymous):

divide 2 from both sides?

jimthompson5910 (jim_thompson5910):

divide both sides by 2, yes

jimthompson5910 (jim_thompson5910):

what do you get when you do this?

OpenStudy (anonymous):

y = -2x +4?

jimthompson5910 (jim_thompson5910):

you have to divide EVERY term by 2, not just the 8

OpenStudy (anonymous):

y=x+4?

jimthompson5910 (jim_thompson5910):

since it works like this \[\Large 2y = -2x + 8\] \[\Large \frac{2y}{2} = \frac{-2x + 8}{2}\] \[\Large \frac{2y}{2} = \frac{-2x}{2} + \frac{8}{2}\] \[\Large y = ??\]

jimthompson5910 (jim_thompson5910):

x+4 is close, but no

OpenStudy (anonymous):

x+3?

jimthompson5910 (jim_thompson5910):

you had the +4 part correct

jimthompson5910 (jim_thompson5910):

the "x" part was incorrect

OpenStudy (anonymous):

why is x incorect?

jimthompson5910 (jim_thompson5910):

you should divide -2x by 2 to get ????

OpenStudy (anonymous):

x?

jimthompson5910 (jim_thompson5910):

nope not x, you are thinking of 2x divided by 2

OpenStudy (anonymous):

isnt that the same thing??

OpenStudy (anonymous):

-x?

jimthompson5910 (jim_thompson5910):

yes -x

jimthompson5910 (jim_thompson5910):

so we will have y = -x + 4

jimthompson5910 (jim_thompson5910):

2x + 2y = 8 turns into y = -x + 4 when you solve for y

jimthompson5910 (jim_thompson5910):

so y = -x + 4 and 2x + 2y = 8 are equivalent (ie the same exact equations)

OpenStudy (anonymous):

thank you for your help

jimthompson5910 (jim_thompson5910):

another term for this is "dependent"

jimthompson5910 (jim_thompson5910):

yw